OpenStudy (anonymous):

What are the real or imaginary solutions of the polynomial equation? 27x3 + 125 = 0 steps too please.

OpenStudy (anonymous):

factor as the sum of two cubes \[a^3+b^3=(a+b)(a^2-ab+b^2)\] use \(a=3x,b=5\)

OpenStudy (anonymous):

then set each factor equal to zero and solve the first answer will be obvious, the second requires the quadratic formula for the two complex zeros

OpenStudy (anonymous):

That's a good formula @satellite73 thanks. I was considering just solving for x and using synthetic division to get rid of the discovered root.